What is the atomic mass (in atomic mass units) of an atom which has a mass of 2.33 × 10-23g?
Answer #1

Answer – We are given , mass of atom = 2.33*10-23 g and we know

1.66054*10-24 g = 1 amu

So, 2.33*10-23 g = ?

= 2.33*10-23 g * 1 amu / 1.66054*10-24 g

= 14.03 amu

So, 14.03 amu is the atomic mass of an atom which has a mass of 2.33 × 10-23g.